kathy borrowed 3300 from the bank in first year of college. how much will she owe at the end of 3 years if her loan is at 6% per year, compounded quarterly?
Answer by mananth(16949)

Principal P = 3300
Amount= A
years=n 3.00
compounded 4 times a year t
Rate = 6.00 0.02
Amount = P*((n+r)/n)^n*t

Amount =3300*( 1+0.02 )^3*4
Amount =3300*(1+0.02)^12
3300 *(1.02)^12
Amount = $ 3945.54